OpenStudy (anonymous):

wat is meant by saying if spring is cutted in pieces then wat will be the relation before cutting spring

OpenStudy (anonymous):

assume the dat spring ius cutted in equal halves

OpenStudy (anonymous):

den wat we can predict abu dat halves?

OpenStudy (wasiqss):

sirf length change hogi cutting mai, spring constant remains same

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

No wasqiss is wrong. If you cut a spring in two then spring constant doubles.

OpenStudy (anonymous):



OpenStudy (anonymous):

mera ans ni a rha isse ques. a spring constant of two springs are K1 and K2 such that one peice is double of the other. then the long peice will have a force constant of??

OpenStudy (wasiqss):

Since f=kx, then k=f/x. Now if you apply the same force to the half-length spring it only moves half as far. So if x is halved above, then k will double. Basically there are now half as many coils to take up the force via displacement.

OpenStudy (anonymous):

well not getting ans dats y i post ques ans is 3/2K

OpenStudy (anonymous):

(3/2)k

OpenStudy (anonymous):

If you cut a spring of constant k into two springs of constants k1 and k2 and lengths l1 and l2 then 1/k=1/k1+ 1/k2. And k1/k2=l2/l1.solve these two equations to get the answer.

OpenStudy (wasiqss):

Let us look at a 10 coil spring as an example. Suppose we apply a force to the spring that compresses it 1 inch meaning that each coil is compressed 1/10 of an inch. If we then cut the spring in half, leaving 5 coils, and then compress it 1 inch, then each coil is compressed 2/10 or 1/5 of an inch. The general spring equation is F = -kx where F is the applied force, k is the spring constant and x is the distance the spring is compressed. So the force needed to compress a single coil by 2/10 of an inch is twice as large as to compress it 1/10 of an inch. Therefore the spring constant (k) of the 5-coil spring must be twice that of the 10-coil spring, because each individual coil of the 5-coil spring is compressed twice as far as each individual coil of the 10-coil spring for the same overall distance. So in general the spring constant of a spring is inversely proportional to the number of coils in the spring.

OpenStudy (anonymous):

guys ans is (3/2)k and shankvee ur formula is rit but can u tell me is K dont changes if l increases plz ??

OpenStudy (anonymous):

K does change if you cut it into two.I don't think i got your question properly....

OpenStudy (anonymous):

ques. a spring constant of two springs are K1 and K2 such that one peice is double of the other. then the long peice will have a force constant of??

OpenStudy (anonymous):

Since one piece is double the other k1/k2=1/2. 1/k=1/k1 +1/k2. So, k1=k2/2, 3/k2=1/k. k2=3k and k1=3/2 k. The longer peice is k1=3/2k

OpenStudy (wasiqss):

yaar heena ques toh sahi likha karo

OpenStudy (anonymous):

sahi toh likha h..

OpenStudy (wasiqss):

shuru mai galat tha

OpenStudy (anonymous):

kaha?? vo to mera doubt tha if u can see

OpenStudy (wasiqss):

ahhh ok dint read that fully

OpenStudy (anonymous):

and shankvee how u get tat springs are arranged in parallel or series??

OpenStudy (anonymous):

toh kisk galti ti... bolo bolo k meri ti

OpenStudy (anonymous):

If two springs have same extension then they are in parallel.

OpenStudy (wasiqss):

ok baba